# Online Planning for Multi-UAV Pursuit-Evasion in Unknown Environments Using Deep Reinforcement Learning

### v3

@@ II RELATED WORK
Reinforcement learning (RL) has emerged as a powerful data-driven approach for solving multi-agent coordination problems in pursuit-evasion games[25, 20, 21, 22, 23, 24]. Several studies have applied RL to enhance traditional methods[25] or address issues such as varying agent numbers[21], efficient communication[20], and credit assignment[29]. However, most RL research remains limited to simplified simulation environments, where agents are modeled as particles or constrained by limited kinematics, reducing the effectiveness of these strategies in real-world scenarios. While recent work has explored deploying RL-based strategies in pursuit-evasion tasks[26, 25, 9, 27], these efforts have largely been confined to two-dimensional environments, such as ground vehicles, surface vehicles or UAVs at fixed altitudes.
@@ IV-A Multi-Agent Reinforcement Learning Setup
Observation Space. The observation $\mathbf{o}_{i}$ for quadrotor $i$ is composed of three components: the self-information $\mathbf{o}_{self}$, states of other quadrotors $\mathbf{o}_{other}$ and information about obstacles $\mathbf{o}_{ob}$. $\mathbf{o}_{self}$ consists of its orientation in quaternions, the linear velocity, the real and predicted future relative position to the evader. If the evader is not detected by any of the quadrotors, we mask the real relative position with a fixed marker value which is $-5$ in our setup. $\mathbf{o}_{other}$ contains the relative positions to other quadrotors. $\mathbf{o}_{ob}$ denotes the relative positions to the $k$-nearest obstacles, where $k$ is smaller than the maximum number of obstacles, representing the quadrotor’s inability to sense global information about environments. In our setup, $k=3$.
@@ IV-B Evader Prediction-Enhanced Network
In the evader prediction-enhanced network, we first build an Evader Prediction Network that predicts the evader’s future trajectories based on historical observations, facilitating the learning of cooperative pursuit strategies under partial observation. The predicted trajectories, concatenated with the raw observations, are then fed into the Actor and Critic Networks. These networks utilize an attention-based observation encoder to better capture the relationships between different entities.
@@ IV-D Two-stage Reward Refinement
We use the smoothness reward $e^{-||a^{t}-a^{t-1}||_{2}}$, where $\boldsymbol{a}_{t}$ represents the policy’s CTBR output, to balance task completion and action smoothness [33]. To address the challenge of maintaining a feasible policy for real-world deployment while ensuring task success, we employ a two-stage reward refinement process. In the first stage, we exclude the smoothness reward and focus on the other three rewards. In the second stage, we reintroduce the smoothness reward and continue training from the first-stage checkpoint. This ensures high task success while improving action smoothness, with the smoothness reward coefficient set to $4.0$.
@@ V-B Evaluation
Evaluation Metrics. We use three statistical metrics to evaluate pursuit strategies: Capture Rate, Capture Step, and Collision Rate. Each experiment is averaged over 300 testing episodes (100 per seed).
@@ V-C Baselines
  • APF [13]: APF guides pursuers to a target position by combining attractive, repulsive, and inter-individual forces. By setting the target of the pursuers to the evader’s position and adjusting the hyperparameters for these forces, the pursuers can navigate towards the evader while avoiding obstacles.
@@
Figure 6: Results of all methods in a obstacle-free scenario. Our algorithm achieves nearly $100\%$ capture rate with all catch radii, and the capture timestep does not increase significantly as the capture radius decreases.
@@ V-E Abaltion Studies
(c) Prediction errors of our method in four test scenarios.
@@
(d) Comparison of our method and variants in OOD scenarios.
Figure 8: Ablation studies. (a, d) Performance of OPEN and baselines on training tasks and OOD generalization. (b) Policy generalizability across evader speeds in the Passage scenario. (c) Evader position prediction error in four test environments.
@@
V-E1 Performance on training tasks:
@@
V-E2 Performance on out-of-distribution scenarios:
We further evaluate our method and its variants in out-of-distribution (OOD) scenarios. As shown in Tab. 8, our method achieves the highest capture rate and the lowest collision rate in the Random scenario, with capture timesteps comparable to“MAPPO + EPN”. In the Passage scenario, our method consistently outperforms all variants. The EPN module provides UAVs with the capability to predict the evader’s strategy, allowing them to plan captures based on the evader’s future trajectory. This significantly reduces the impact of OOD scenarios on policy performance. Additionally, the AEG module enhances the policy’s ability to address corner cases, promoting better generalization across diverse environments. By integrating these two modules, our method surpasses all baseline variants.
@@ V-F Real-World Experiments
Scenarios Wall Narrow Gap Random Passage
Sim. 200.6(01.0) 348.6(16.2) 292.4(00.9) 361.0(00.4)
Real. 320.2(26.8) 441.7(82.1) 385.2(29.4) 377.5(49.2)
TABLE II: Capture steps for two-stage reward refinement in four test scenarios, evaluated in both simulation and real-world settings. The results demonstrate that the two-stage reward refinement achieves strategies in the real world that closely mirror those observed in simulation.
As shown in Fig. 9, we deploy the policy on three real CrazyFlie 2.1 quadrotors, each with a maximum speed of 1.0 m/s. A motion capture system provides the quadrotors’ states, while a virtual evader’s true states are input into the actor and evader prediction network upon detection. The actor and evader prediction network run on a local computer, sending CTBR control commands at 100 Hz to the quadrotors via radio.
Figure 9: Snapshot of the real-world experiments.
We compare two-stage reward refinement (Two-stage) with its variants: training without the smoothness reward (w/o Smoothness) and direct training using all rewards, including the smoothness reward (One-stage). In real-world experiments, w/o Smoothness fails entirely, as drastic output command changes cause the drone to crash. One-stage, using the same smoothness reward coefficients as Two-stage, achieves a capture success rate of 0 in simulation. This is because the smoothness reward minimizes output command variations, conflicting with capture objectives and causing UAV inactivity. In contrast, Two-stage is the only method enabling stable real-world deployment. As shown in Tab. II, we compare capture steps between simulation and real-world environments using the same initial positions, repeating the test five times across four scenarios. The number of capture steps in the real world is slightly higher than in simulation, but the strategy remains consistent with the simulation. A real-world demonstration video is provided in the supplementary materials.
@@ VI CONCLUSION
We learn a feasible RL policy that can perform online planning in the unknown environment for multi-UAV pursuit- evasion and deploy it on real quadrotors. We introduce an Adaptive Environment Generator to automatically generate curricula for policy generalization across diverse scenarios and employ an Evader Prediction-Enhanced Network for cooperatively capture. Simulations show our method significantly outperforms all baselines and generalizes to unseen scenarios. We employ a two-stage reward refinement process to ensure the policy adheres to physical constraints and deploy it on real Crazyflie quadrotors in a zero-shot manner.
Although we first derive and deploy an RL-based policy using CTBR commands for multi-UAV pursuit-evasion in unknown environments, several limitations remain. For example, although the policy supports varying numbers of drones, it still requires retraining to ensure performance when the number of drones changes. Additionally, the current modeling of obstacles is only suitable for structured environments, and future work should explore vision-based pursuit-evasion tasks to extend this approach to unstructured environments.

### v4

@@ Abstract
Index Terms:
Reinforcement Learning; Cooperating Robots; Multi-Robot Systems
@@ II RELATED WORK
Reinforcement learning (RL) has emerged as a powerful data-driven approach for solving multi-agent coordination problems in pursuit-evasion games [25, 20, 21, 22, 23, 24, 29]. Several studies have applied RL to enhance traditional methods [25] or address issues such as varying agent numbers [21], efficient communication [20], credit assignment [30] or perception uncertainty [29]. However, most RL research remains limited to simplified simulation environments, where agents are modeled as particles or constrained by limited kinematics, reducing the effectiveness of these strategies in real-world scenarios. While recent work has explored deploying RL-based strategies in pursuit-evasion tasks [26, 25, 9, 27], these efforts have largely been confined to two-dimensional environments, such as ground vehicles, surface vehicles or UAVs at fixed altitudes.
@@ IV-A Multi-Agent Reinforcement Learning Setup
Observation Space. The observation $\mathbf{o}_{i}$ for quadrotor $i$ is composed of three components: the self-information $\mathbf{o}_{self}$, states of other quadrotors $\mathbf{o}_{other}$ and information about obstacles $\mathbf{o}_{ob}$. $\mathbf{o}_{self}$ consists of its orientation in quaternions, the linear velocity, the real and predicted future relative position to the evader. If the evader is not detected by any of the quadrotors, we mask the real relative position with a fixed marker value which is $-5$ in our setup. $\mathbf{o}_{other}$ contains the relative positions to other quadrotors. $\mathbf{o}_{ob}$ denotes the relative positions to the $k$-nearest obstacles, where $k$ is smaller than the maximum number of obstacles, representing the quadrotor’s inability to sense global information. In our setup, $k=3$.
@@ IV-B Evader Prediction-Enhanced Network
In the evader prediction-enhanced network, we build an Evader Prediction Network that predicts the evader’s future trajectories based on historical observations, facilitating the learning of cooperative pursuit strategies under partial observation. The predicted trajectories, concatenated with the raw observations, are then fed into the Actor and Critic Networks. These networks utilize an attention-based observation encoder to better capture the relationships between different entities.
@@ IV-D Two-stage Reward Refinement
We use the smoothness reward $e^{-||a^{t}-a^{t-1}||_{2}}$, where $\boldsymbol{a}_{t}$ represents the policy’s CTBR output, to balance task completion and action smoothness [34]. To address the challenge of maintaining a feasible policy for real-world deployment while ensuring task success, we employ a two-stage reward refinement process. In the first stage, we exclude the smoothness reward and focus on the other three rewards. In the second stage, we reintroduce the smoothness reward and continue training from the first-stage checkpoint. This ensures high task success while improving action smoothness, with the smoothness reward coefficient set to $2.0$.
@@ V-B Evaluation
Evaluation Metrics. We use three statistical metrics to evaluate pursuit strategies: Capture Rate, Capture Step, and Collision Rate. Each experiment is averaged over 300 testing episodes (100 per seed). All numerical values are reported as mean (standard deviation).
@@ V-C Baselines
  • APF [13]: APF guides pursuers to a target position by combining attractive, repulsive, and inter-individual forces.
@@
Figure 6: Results of all methods in a obstacle-free scenario.
@@ V-E Ablation Studies
Figure 8: Ablation studies. (a) Performance of OPEN and baselines on training tasks. (b) Policy generalization across evader speeds in the Passage scenario.
@@
TABLE II: Prediction errors of our method in four test scenarios.
@@
TABLE III: Comparison of OPEN and variants in OOD scenarios.
@@
V-E1 Performance on training tasks
@@
V-E2 Performance on out-of-distribution scenarios
We further evaluate our method and its variants in out-of-distribution scenarios. As shown in Tab. III, our method achieves the highest capture rate and the lowest collision rate in the Random scenario, with capture timesteps comparable to“MAPPO + EPN”. In the Passage scenario, our method consistently outperforms all variants. The EPN module provides UAVs with the capability to predict the evader’s strategy, allowing them to plan captures based on the evader’s future trajectory. This significantly reduces the impact of OOD scenarios on policy performance. Additionally, the AEG module enhances the policy’s ability to address corner cases, promoting better generalization across diverse environments. By integrating these two modules, our method surpasses all baseline variants.
@@ V-F Real-World Experiments
Methods Envs. Wall Narrow Gap Random Passage
w/o Smoothness (oracle) Sim. 258.6(0.00) 478.0(0.00) 361.9(0.00) 362.0(0.00)
w/o Smoothness Real. fail fail fail fail
One-stage Real. / / / /
Two-stage Real. 320.2(26.8) 441.7(82.1) 385.2(29.4) 377.5(49.2)
TABLE IV: Capture steps of all methods in four test scenarios. "/" notes that the One-stage fails to generate effective pursuit strategies in the simulation and we do not conduct experiments in the real-world.
We deploy the policy on three real CrazyFlie 2.1 quadrotors, each with a maximum speed of 1.0 m/s. A motion capture system provides the quadrotors’ states, while a virtual evader’s true states are input into the actor and evader prediction network upon detection. The actor and evader prediction network run on a local computer, sending CTBR control commands at 100 Hz to the quadrotors via radio. We align the initial conditions of the simulation with real-world to ensure fair comparison. All real experiments are repeated 5 times and recorded as the mean and standard deviation.
We compare the two-stage reward refinement (Two-stage) with its variants. We define the first stage of Two-stage as w/o Smoothness, training only with task rewards. So the simulation results of w/o Smoothness can be considered as oracle. One-stage denotes training with both smoothness and task rewards. As shown in Tab. IV, the results demonstrate that w/o Smoothness fails completely in real-world experiments due to overly aggressive pursuit strategies which lead to unstable behaviors and UAV crashes. One-stage causes severe exploration challenges in MARL training, resulting in no policy improvement during training and yielding only trivial stationary strategies. In contrast, only the Two-stage achieves successful real-world deployment while preserving capture performance approximately consistent with simulation results.
@@ VI CONCLUSION
We propose a feasible RL policy for online planning in unknown environments, enabling real-world deployment on Crazyflie quadrotors for multi-UAV pursuit-evasion. To promote generalization, we introduce an Adaptive Environment Generator for automatic curriculum generation and a Prediction-Enhanced Network for cooperative capture. Our method outperforms baselines in simulation and generalizes well to unseen scenarios. A two-stage reward refinement ensures physical feasibility and enables zero-shot real-world transfer.
Our method still has some limitations. The policy supports varying drone numbers but still needs retraining for performance. Moreover, future work will explore vision-based methods for unstructured environments.